princeevee:

A group of people in an art class and a crafts class were surveyed whether they exercise more than 5 hours a week or less than 5 hours a week. In the art class, 5 people exercise more than 5 hours a week and 8 people exercise less than 5 hours a week. In the crafts class, 10 people exercise more than 5 hours a week and 4 people exercise less than 5 hours a week. What is the relative frequency of people who exercise more than 5 hours a week in the art class to the total number of people surveyed? Round the answer to the nearest hundredth.

ThatCreepyBoy:

I got this from os ~ "

ThatCreepyBoy:

ftw...."

ThatCreepyBoy:

The Relative Frequency is how often something happens out of all possible outcomes. Here you have numbers -- Arts - 5 more, 8 less Crafts- 10 more, 4 less


What is the relative frequency of people who exercise more than 5 hours a week in Arts class? There are 5 people in arts that exercise more than 5 hours/week. The total number of people all together surveyed is 5 + 8 + 10 + 4 = 27 people There are 5 out of 27 people that are in this catagory Relative Frequency = 5/27 = 0.19

ThatCreepyBoy:

The Relative Frequency is how often something happens out of all possible outcomes. Here you have numbers -- Arts - 5 more, 8 less Crafts- 10 more, 4 less


What is the relative frequency of people who exercise more than 5 hours a week in Arts class? There are 5 people in arts that exercise more than 5 hours/week. The total number of people all together surveyed is 5 + 8 + 10 + 4 = 27 people There are 5 out of 27 people that are in this catagory Relative Frequency = 5/27 = 0.19

princeevee:

thanks

ThatCreepyBoy:

Np

princeevee:

1 attachment
princeevee:

can you help with this?

ThatCreepyBoy:

Sure give me one second :D

ThatCreepyBoy:

Dang o.o nooooo d,x @Vocaloid Can ya help?

Vocaloid:

hint: since both variables are increasing, would this be positive or negative?

Vocaloid:

 is this positive or negative?

ThatCreepyBoy:

@princeevee

princeevee:

positive

princeevee:

i assume

Vocaloid:

good job! so positive is your answer

ThatCreepyBoy:

Good job voca ^_^